OpenStudy (anonymous):

Let f(t) = t2 + 3 and g(t) = 4t - 3 Find (f o g)(-1)

OpenStudy (anonymous):

g(-1) = 4(-1) - 3 = -7 Plug that into f function, f(-7) = (-7)^2 +3 = 49 +3 = 52 I assume t2 meant t squared?

OpenStudy (anonymous):

yes :) && thank you SO much <3